# Accurate Screening of Functional Materials with Machine-Learning Potential and Transfer-Learned Regressions: Heusler Alloy Benchmark

### v1

@@ 2.2 Distribution of strong $E_{\mathrm{aniso}}$ candidates
Figure 3: Distribution of ML-selected candidate compounds based on whether 4d or 5d elements are present and distribution over 4d and 5d elements contained. The distribution of DFT validated strong $E_{\mathrm{aniso}}$ candidates is also shown.
@@
2.3 MLIPs optimization performance
@@
Figure 4: Lattice constants $a$ and $c$, $c/a$ ratio, total energy ($E$), formation energy ($\Delta E$), and convex hull distance ($\Delta H$) predicted by various MLIPs are benchmarked against DFT references. For each property, the fraction of compounds with predictions falling within specified relative error (RE) or absolute error (AE) thresholds is reported. Energetic quantities ($E$, $\Delta E$, and $\Delta H$) are expressed in eV/atom.
@@
In recent years, MLIPs have advanced rapidly, with numerous new models proposed and trained. To identify the most suitable model for our screening workflow, we benchmarked representative MLIPs models developed since 2023. These include ALIGNN-FF, CHGNet, SevenNet-l3i5, SevenNet-MF-ompa, HIENet, MatterSim-v1, eqV2-S-OAM, eqV2-M-OAM, eqV2-L-OAM, and the latest eSEN-30M-OAM [22, 23, 24, 25, 26, 13, 21]. The evaluation focused on structure optimization for 10,000 conventional ternary compounds randomly selected from the ground states in HeuslerDB. To identify global minimum, 14 initial structures were generated by applying strain to the conventional cell (two formula units) and converting it to the primitive cell (one formula unit). Specifically, the $a$, $b$, and $c$ axes were uniformly scaled by $\pm 10\%$ and $\pm 30\%$, or the $c$-axis alone was varied by $\pm 10\%$, $\pm 20\%$, $\pm 30\%$, $\pm 40\%$, and $\pm 50\%$.
@@
The performance of structure optimization was assessed by comparing the predicted lattice constants $a$ and $c$, and the resulting $c/a$ ratio, with corresponding DFT values. The results are summarized in Fig. 4. The relative error (RE) is defined as the maximum of $\left|\frac{x_{\mathrm{ML}}}{x_{\mathrm{DFT}}}-1\right|$ and $\left|\frac{x_{\mathrm{DFT}}}{x_{\mathrm{ML}}}-1\right|$, where $x_{\mathrm{ML}}$ and $x_{\mathrm{DFT}}$ denote the values predicted by the ML and DFT, respectively. We report the fractions of compounds within 5% and 1% RE tolerances. Among the evaluated models, the eSEN-30M-OAM and eqV2 models achieved the highest accuracy at 5% RE, with eSEN-30M-OAM showing slightly better performance at the stricter 1% RE threshold. A key distinction between the two models lies in the number of local minima encountered: eqV2-L-OAM (91,585) produced substantially more than eSEN-30M-OAM (32,606). This difference arises from the smoother energy landscape of eSEN-30M-OAM, enabling fewer initial structures to converge to the global minimum and thereby improving optimization efficiency. Convergence tests of all evaluated models are provided in the Supplementary Materials.
@@
The predictive performance of total energy ($E$), formation energy ($\Delta E$), and energy above the convex hull ($\Delta H$) using MLIP was assessed by comparing the MLIP predictions with DFT values, using the absolute error (AE), $|x_{\mathrm{ML}}-x_{\mathrm{DFT}}|$. The fractions of compounds with AE below 0.01 and 0.05 eV/atom are shown in Fig. 4. Among the benchmarked models, eSEN-30M-OAM and the eqV2 variants showed the highest accuracy for $\Delta E$ and $\Delta H$ at the 0.05 eV/atom threshold, with eSEN-30M-OAM displaying a slight drop in accuracy at the more stringent 0.01 eV/atom level. Predicted total energies from MLIP were found to be systematically lower than DFT values, reducing direct agreement; however, this offset also applies to elemental references and competing phases, so the relative quantities $\Delta E$ and $\Delta H$ remain in strong agreement with DFT. Given its robust performance in both structure optimization and thermodynamic stability, eSEN-30M-OAM was selected for integration into the ML-HTP workflow.
@@
Previous studies have reported the performance of composition-based models in predicting lattice constants, $\Delta E$, and $\Delta H$ for cubic Heusler compounds. For comparison, we evaluated metric of eSEN-30M-OAM on the cubic Heusler subset, with results summarized in Table 1. The $R^{2}$ score for the lattice constant $a$ is 0.994, surpassing the previously reported ranges of 0.80–0.94 across different Heusler types and the values of 0.94, 0.979, and 0.987 in other works [8, 27, 28, 10]. Similarly, the $R^{2}$ for $\Delta E$ reaches 0.995, outperforming earlier results of 0.80–0.88 and 0.93, 0.982 [8, 27, 10]. The $R^{2}$ for $\Delta H$ is 0.98, exceeding prior values of 0.91 and 0.969 [29, 10]. The root mean squared errors (RMSE) for $a$ and $\Delta E$ are 0.02 Å and 0.029eV/atom, respectively, which are significantly lower than the 0.11Å and 0.117eV/atom reported in earlier work[9].
@@
Table 1: Performance comparison of the eSEN MLIP and MLRM with ALIGNN models and previously reported results. The size of dataset used for MLRM is also listed.
@@
$\{\boldsymbol{m}_{i}\}$ $R^{2}$ 0.990 27,864
$m_{\mathrm{total}}$ $R^{2}$ 0.986 0.90 0.75–0.8911footnotemark: 1, 0.8277footnotemark: 7, 0.9033footnotemark: 3 27,864
$\sum|\mathbf{m}_{i}|$ $R^{2}$ 0.990 0.89 27,864
$\omega_{\mathrm{min}}$ $R^{2}$ 0.750 0.73 8,198
$T_{{\rm c}}$ $R^{2}$ 0.910 0.84 0.7688footnotemark: 8, 0.7399footnotemark: 9 2,106
@@
00footnotetext: 1[8] Dataset size is about 1000, 2[27] Dataset size is about 65,000,
3[10] Dataset size is about 500,000 for $a$ and $m_{\mathrm{total}}$, and about 450,000 for $\Delta E$ and $\Delta H$,
4[28] Dataset size is 143, 5[9] Dataset size is 16,272, 6[29] Dataset size is 426,148,
7[30] Dataset size is 1153, 8[6] Dataset size is 408, 9[31] Dataset size is 6500.
2.4 Improvements Over Existing Approaches
The MLRM s used in screening were trained on the data from HeuslerDB, supplemented with newly computed $T_{{\rm c}}$ and $E_{\mathrm{aniso}}$ values using optimized structures in HeuslerDB. Test set metrics are summarized in Table 1 and benchmarked against previously reported results. The MLRM for $\{\boldsymbol{m}_{i}\}$ achieved an $R^{2}$ score of 0.990. For comparison with prior studies that used total magnetization ($m_{\mathrm{total}}$) as the target property, our model yielded $R^{2}=0.986$ for $m_{\mathrm{total}}$, exceeding earlier values of 0.75–0.89, 0.82, and 0.90 [8, 30, 10]. For $T_{{\rm c}}$, the model attained $R^{2}=0.91$ and classification accuracy of 0.91, both substantially higher than previously reported values of $R^{2}=0.76$ and 0.73 and accuracy of 0.73 [6, 31].
To the best of our knowledge, no previous study has predicted phonon stability or $E_{\mathrm{aniso}}$ of Heusler compounds using ML. The effectiveness of the $\omega_{\mathrm{min}}$ and $E_{\mathrm{aniso}}$ models developed here is supported by the validation results in Sec. 2. While phonon stability could also be assessed using MLIP combined with conventional phonon calculation methods, a detailed comparison between this MLIP-based strategy and the MLRM approach adopted in our workflow is left for future work focused on phonon properties and their interactions. The $E_{\mathrm{aniso}}$ model achieved an $R^{2}$ of 0.68—lower than those of $\{\boldsymbol{m}_{i}\}$, $\omega_{\mathrm{min}}$, and $T_{{\rm c}}$ —highlighting the higher sensitivity and complexity of $E_{\mathrm{aniso}}$ as a target property. Nevertheless, despite the reduced $R^{2}$, its classification accuracy remains satisfactory and sufficient for integration into the ML-HTP workflow.
@@ 2.4 Improvements Over Existing Approaches
We further tested a hybrid workflow in which structure optimization was performed with eSEN-30M-OAM MLIP, while property prediction was carried out using ALIGNN MLRMs. This approach yielded 276 candidate compounds, of which 149 (54.0%) were confirmed by DFT to exhibit strong $E_{\mathrm{aniso}}$. The improved yield relative to ALIGNN-FF based optimization underscores the critical importance of accurate structure optimization with eSEN-30M-OAM for enhancing ML-HTP screening. However, the yield still falls short of the 82.0% achieved by the fully eSEN-based workflow, indicating that progress in both the MLIP and MLRM components is essential for maximal efficiency. We also tested the inverse hybrid configuration, using ALIGNN-FF MLIP for structure optimization combined with eSEN MLRMs for property prediction. This workflow identified 243 candidates, of which only 76 (31.3%) were validated as strong $E_{\mathrm{aniso}}$ compounds. This marked reduction in performance highlights the pivotal role of selecting an accurate MLIP for structure optimization.
@@ 2.5 Prediction of local magnetic moment
Since the goal of this study is to identify compounds with large $E_{\mathrm{aniso}}$, it is first necessary to determine whether a compound is magnetic. Relying solely on the total magnetization is inadequate, as it cannot capture antiferromagnetic (AFM) or low-moment ferrimagnetic (FiM) compounds. To address this, we employed the total absolute magnetic moment, defined as $\sum_{i}|\boldsymbol{m}_{i}|$, where $\boldsymbol{m}_{i}$ denotes the local magnetic moment at atomic site $i$.
Local magnetic moments were predicted using an MLRM based on the eSEN architecture, trained to output the moment at each atomic site. Restricting to collinear configurations that all moments aligned along the $z$-axis, each moment is represented by a scalar whose sign encodes the direction, with $\ell=0$ per site in the output head. To account for the $z$-direction ambiguity—where a configuration and its sign-inverted counterpart (i.e., all local moments flipped) are physically equivalent—we modified the loss function to compute losses for both the predicted $\{\boldsymbol{m}_{i}\}$ and its sign-inverted counterpart, and use the smaller value as the loss. This ensures invariance under global spin inversion.
Fig 5 (a) shows a scatter plot comparing $\{\boldsymbol{m}_{i}\}$ from the MLRM and DFT for compounds in test set, with histograms along the axes illustrating their distributions. Local moments at the atom site $Z$ are omitted as they are non-magnetic. Because 73.4% of the test compounds are non-magnetic, both histograms exhibit a pronounced peak at zero. For magnetic systems, the global sign is adjusted so that the total magnetic moment is positive; since most magnetic compounds are ferromagnetic, positive moments dominate in the distribution. Nearly all points fall along the diagonal, demonstrating that the model accurately reproduces both the magnitude and sign of local moments, and reliably distinguishes ferromagnetic (FM) and ferrimagnetic (FiM) systems. Only 1.4% of points lie along either axis, corresponding to DFT-magnetic sites incorrectly predicted as non-magnetic or vice versa, indicating highly accurate magnetic/non-magnetic classification.
Predicting $\{\boldsymbol{m}_{i}\}$ is a common yet computationally demanding step in HTP studies of magnetic materials. The approach developed in this work achieves high accuracy in $\{\boldsymbol{m}_{i}\}$ prediction and is readily transferable to other systems. A central question, however, is how many training compounds are required to reach satisfactory accuracy. To address this, we performed a learning-curve analysis by training the MLRM on progressively larger subsets of the dataset and evaluating performance on a fixed validation set of 5000 compounds, of which 1486 are magnetic. For evaluation, local moments at the $X_{1}$, $X_{2}$, and $Y$ sites were concatenated across samples into a single array, while the $Z$ site was excluded since it is non-magnetic in conventional Heusler compounds.
The learning curve is shown in Fig. 5 (b), illustrating how model performance improves with increasing training set size. In the top panel, the $R^{2}$ scores for both local moments and their magnitudes are presented. The gap between the two curves indicates that, while the model generally captures the magnitude accurately, it sometimes assigns the incorrect sign. For example, Mn2ScGe with DFT-computed local moments {$2.62$, $3.03$, 0.29-0.29, 0.10-0.10} $\mu_{\mathrm{B}}$ is predicted as {2.71-2.71, $3.04$, $0.02$, 0.02-0.02} $\mu_{\mathrm{B}}$ by MLRM. The bottom panel reports two key metrics: (i) classification accuracy for identifying magnetic systems and (ii) the fraction of local moments with absolute error below 0.1 $\mu_{\mathrm{B}}$, evaluated across all compounds and within the magnetic subset. With 5000 training samples, the model achieves a classification accuracy of 0.92, and 90% of all compounds fall within the 0.1 $\mu_{\mathrm{B}}$ error threshold. However, this fraction decreases to 0.72 when restricted to magnetic compounds, indicating that the model identifies whether a site is magnetic with high reliability but remains less accurate in predicting exact $\{\boldsymbol{m}_{i}\}$ values. Increasing the training set to 125,000 samples improves this fraction to 0.82, while relaxing the threshold to 0.2 $\mu_{\mathrm{B}}$ further raises it to 0.92. Although performance improves with larger datasets, the gains become progressively smaller. These results highlight the critical role of dataset size in improving $\{\boldsymbol{m}_{i}\}$ accuracy and inform the selection of training size in future work focusing on other magnetic systems.
@@
Figure 5: (a) Scatter plot comparing ML-predicted $\{\boldsymbol{m}_{i}\}$ with DFT values for the test set. (b) Learning curves for $\{\boldsymbol{m}_{i}\}$ prediction. The top panel shows $R^{2}$ scores for both local moments and their magnitudes. The bottom panel reports the fraction of compounds with absolute prediction error below 0.1 $\mu_{\mathrm{B}}$ for all compounds and for the magnetic subset, together with magnetic system classification accuracy. (c) $R^{2}$ of models initialized from the eSEN-30M-OAM MLIP and trained with different numbers of frozen layers (denoted TL-MLIP-$n$). In the $n=0$ case, the embedding layer is also trainable. A model trained from scratch (denoted w/o TL) and a transfer learning variant initialized from a base model pre-trained on $\Delta E$ data from the HeuslerDB database (denoted TL-$\Delta E$-0) are included for comparison.

### v2

@@ 2.2 Distribution of strong $E_{\mathrm{aniso}}$ candidates
Figure 3: Distribution of ML-selected compounds on elements contained. Distribution of ML-selected candidate compounds based on whether 4d or 5d elements are present and distribution over 4d and 5d elements contained. The distribution of DFT validated strong $E_{\mathrm{aniso}}$ candidates is also shown.
@@
2.3 uMLIP optimization performance
@@
Figure 4: Benchmark of uMLIP performance. Lattice constants $a$ and $c$, $c/a$ ratio, total energy ($E$), formation energy ($\Delta E$), and convex hull distance ($\Delta H$) predicted by various uMLIPs are benchmarked against DFT references. For each property, the fraction of compounds with predictions falling within specified relative error (RE) or absolute error (AE) thresholds is reported. Energetic quantities ($E$, $\Delta E$, and $\Delta H$) are expressed in eV/atom. The test set consists of 10,000 ground-state compounds randomly sampled from HeuslerDB.
@@
In recent years, uMLIPs have advanced rapidly, with numerous new models proposed and trained. To identify the most suitable model for our screening workflow, we benchmarked representative uMLIPs models developed since 2023. These include ALIGNN-FF, CHGNet, SevenNet-l3i5, SevenNet-MF-ompa, HIENet, MatterSim-v1, eqV2-S-OAM, eqV2-M-OAM, eqV2-L-OAM, and the latest eSEN-30M-OAM [choudharyUnifiedGraphNeural2023, dengCHGNetPretrainedUniversal2023, kimDataEfficientMultifidelityTraining2025, yanMaterialsFoundationModel2025, yangMatterSimDeepLearning2024, barroso-luqueOpenMaterials20242024, fuLearningSmoothExpressive2025]. The evaluation focused on structure optimization for 10,000 conventional ternary compounds randomly selected from the ground states in HeuslerDB. To identify the global minimum, 14 initial structures were generated by applying strain to the conventional cell (two formula units) and converting it to the primitive cell (one formula unit). Specifically, the $a$, $b$, and $c$ axes were uniformly scaled by $\pm 10\%$ and $\pm 30\%$, or the $c$-axis alone was varied by $\pm 10\%$, $\pm 20\%$, $\pm 30\%$, $\pm 40\%$, and $\pm 50\%$. The lowest-energy structure from these relaxations was selected as the predicted ground state. Convergence tests for all evaluated models are shown in Fig. S4, demonstrating that the selected ground states are well-converged.
@@
The performance of structure optimization was assessed by comparing the predicted lattice constants $a$ and $c$, and the resulting $c/a$ ratio, with corresponding DFT values. The results are summarized in Fig. 4. The relative error (RE) is defined as the maximum of $\left|\frac{x_{\mathrm{ML}}}{x_{\mathrm{DFT}}}-1\right|$ and $\left|\frac{x_{\mathrm{DFT}}}{x_{\mathrm{ML}}}-1\right|$, where $x_{\mathrm{ML}}$ and $x_{\mathrm{DFT}}$ denote the values predicted by the ML and DFT, respectively. We report the fractions of compounds within 5% and 1% RE tolerances. Among the evaluated models, the eSEN-30M-OAM and eqV2 models achieved the highest accuracy at 5% RE, with eSEN-30M-OAM showing slightly better performance at the stricter 1% RE threshold. A key distinction between the two models is the number of local minima encountered: eqV2-L-OAM identified 91,585 local minima, whereas eSEN-30M-OAM identified 32,606. Both counts are much greater than 10,000, confirming the presence of multiple local minima, but are still significantly less than 10,000×14. This suggests that many different initial distortions ultimately converge to the same local minimum. Importantly, eSEN-30M-OAM found substantially fewer local minima than eqV2-L-OAM. This difference is attributed to the smoother energy landscape of eSEN-30M-OAM. Convergence tests also demonstrate that eSEN-30M-OAM achieves convergence with fewer initial structures, resulting in lower computational cost for HTP purpose.
@@
The predictive performance of total energy ($E$), formation energy ($\Delta E$), and energy above the convex hull ($\Delta H$) using uMLIP was assessed by comparing the uMLIP predictions with DFT values, using the absolute error (AE), $|x_{\mathrm{ML}}-x_{\mathrm{DFT}}|$. The fractions of compounds with AE below 0.01 and 0.05 eV/atom are shown in Fig. 4. Among the benchmarked models, eSEN-30M-OAM and the eqV2 variants showed the highest accuracy for $\Delta E$ and $\Delta H$ at the 0.05 eV/atom threshold, with eSEN-30M-OAM displaying a slight drop in accuracy at the more stringent 0.01 eV/atom level. Predicted total energies from uMLIP were found to be systematically lower than DFT values, reducing direct agreement; however, this offset also applies to elemental references and competing phases, so the relative quantities $\Delta E$ and $\Delta H$ remain in strong agreement with DFT. Given its robust performance in both structure optimization and thermodynamic stability, eSEN-30M-OAM was selected for integration into the ML-HTP workflow.
2.4 Improvements Over Existing Approaches
@@
Previous studies have reported the performance of composition-based models in predicting lattice constants, $\Delta E$, and $\Delta H$ for cubic Heusler compounds. For comparison, we evaluated metrics of eSEN-30M-OAM on the cubic Heusler subset, with results summarized in Table 1. The $R^{2}$ score for the lattice constant $a$ is 0.994, surpassing the previously reported ranges of 0.80–0.94 across different Heusler types and the values of 0.94, 0.979, and 0.987 in other works [mitraMachineLearningApproach2023, huSearchingHighSpin2020, miyazakiMachineLearningBased2021, luExplainableAttentionCNN2025]. Similarly, the $R^{2}$ for $\Delta E$ reaches 0.995, outperforming earlier results of 0.80–0.88 and 0.93, 0.982 [mitraMachineLearningApproach2023, huSearchingHighSpin2020, luExplainableAttentionCNN2025]. The $R^{2}$ for $\Delta H$ is 0.98, exceeding prior values of 0.91 and 0.969 [kimMachinelearningacceleratedHighthroughputMaterials2018, luExplainableAttentionCNN2025]. The root mean squared errors (RMSE) for $a$ and $\Delta E$ are 0.023 Å and 0.029 eV/atom, respectively, which are significantly lower than the 0.11-0.12 Å and 0.117 eV/atom reported in earlier work [xieScreeningNewQuaternary2023].
The MLRM s used in screening were trained on the data from HeuslerDB, supplemented with newly computed $T_{{\rm c}}$ and $E_{\mathrm{aniso}}$ values using optimized structures in HeuslerDB. Test set metrics are summarized in Table 1 and benchmarked against previously reported results. The MLRM for $\{\boldsymbol{m}_{i}\}$ achieved an $R^{2}$ score of 0.989. For comparison with prior studies that used total magnetization ($m_{\mathrm{total}}$) as the target property, our model yielded $R^{2}=0.986$ for $m_{\mathrm{total}}$, exceeding earlier values of 0.75–0.89, 0.82, and 0.927 [mitraMachineLearningApproach2023, liuMachineLearningAssisted2025, luExplainableAttentionCNN2025]. For $T_{{\rm c}}$, the model attained $R^{2}=0.91$ and classification accuracy of 0.91, both substantially higher than previously reported values of $R^{2}=0.76$ and 0.73 and accuracy of 0.73 [hilgersMachineLearningbasedEstimation2025, hirohataMachineLearningDevelopment2025].
@@
Table 1: Performance comparison of the eSEN uMLIP and MLRM with ALIGNN models and previously reported results. The size of dataset used for MLRM is also listed.00footnotetext: 1[mitraMachineLearningApproach2023] Dataset size is about 1000, 2[huSearchingHighSpin2020] Dataset size is about 65,000,
3[luExplainableAttentionCNN2025] Dataset size is about 500,000 for $a$ and $m_{\mathrm{total}}$, and about 450,000 for $\Delta E$ and $\Delta H$,
4[miyazakiMachineLearningBased2021] Dataset size is 143, 5[xieScreeningNewQuaternary2023] Dataset size is 16,272, 6[kimMachinelearningacceleratedHighthroughputMaterials2018] Dataset size is 426,148,
7[liuMachineLearningAssisted2025] Dataset size is 1153, 8[hilgersMachineLearningbasedEstimation2025] Dataset size is 408, 9[hirohataMachineLearningDevelopment2025] Dataset size is 6500.
@@
$\{\boldsymbol{m}_{i}\}$ $R^{2}$ 0.989 27,864
$m_{\mathrm{total}}$ $R^{2}$ 0.986 0.904 0.75–0.8911footnotemark: 1, 0.8277footnotemark: 7, 0.92733footnotemark: 3 27,864
$\sum|\mathbf{m}_{i}|$ $R^{2}$ 0.989 0.891 27,864
$\omega_{\mathrm{min}}$ $R^{2}$ 0.750 0.734 8,198
$T_{{\rm c}}$ $R^{2}$ 0.910 0.844 0.7688footnotemark: 8, 0.7399footnotemark: 9 2,106
@@
To the best of our knowledge, no previous study has predicted phonon stability or $E_{\mathrm{aniso}}$ of Heusler compounds using ML. The effectiveness of the $\omega_{\mathrm{min}}$ and $E_{\mathrm{aniso}}$ models developed here is supported by the validation results in Sec. 2. While phonon stability could also be assessed using uMLIP combined with phonon calculation methods, our regressor-based approach is motivated by both efficiency and accuracy. Only the minimum frequency is needed for phonon stability assessment, thus a regressor is sufficient and much faster than calculating the full spectrum. We also evaluated uMLIP +phonon using CHGNet, MatterSim, and eSEN-30M-OAM on a test set of 1000 conventional ternary compounds, and found stability/instability classification accuracies of 62.5%, 74.9%, and 80.2%, respectively, which are substantially below the regressor’s 93.6%. Notably, CHGNet misclassifies 67.2% of stable compounds as unstable, showing a strong tendency to underestimate stability, while eSEN weakly overestimates it and MatterSim exhibits a more balanced performance. Additional details including example phonon spectra by uMLIP s and DFT are in the Supplementary Information. The $E_{\mathrm{aniso}}$ model achieved an $R^{2}$ of 0.68, lower than those of $\{\boldsymbol{m}_{i}\}$, $\omega_{\mathrm{min}}$, and $T_{{\rm c}}$, highlighting the higher sensitivity and complexity of $E_{\mathrm{aniso}}$ as a target property. Nevertheless, despite the reduced $R^{2}$, its classification accuracy remains satisfactory and sufficient for integration into the ML-HTP workflow.
@@
We further tested a hybrid workflow in which structure optimization was performed with eSEN-30M-OAM uMLIP, while property prediction was carried out using ALIGNN MLRMs. This approach yielded 276 candidate compounds, of which 149 (54.0%) were confirmed by DFT to exhibit strong $E_{\mathrm{aniso}}$. The improved yield relative to ALIGNN-FF based optimization underscores the critical importance of accurate structure optimization with eSEN-30M-OAM for enhancing ML-HTP screening. However, the yield still falls short of the 82.0% achieved by the fully eSEN-based workflow, indicating that progress in both the uMLIP and MLRM components is essential for maximal efficiency. We also tested the inverse hybrid configuration, using ALIGNN-FF uMLIP for structure optimization combined with eSEN MLRMs for property prediction. This workflow identified 243 candidates, of which only 76 (31.3%) were validated as strong $E_{\mathrm{aniso}}$ compounds. This marked reduction in performance highlights the pivotal role of selecting an accurate uMLIP for structure optimization.
@@ 2.5 Prediction of local magnetic moment
Figure 5: Local magnetic moment prediction performance. a) Scatter plot comparing ML-predicted $\{\boldsymbol{m}_{i}\}$ with DFT values for the test set. b) Learning curves for $\{\boldsymbol{m}_{i}\}$ prediction. The top panel shows $R^{2}$ scores for both local moments and their magnitudes. The bottom panel reports the magnetic/nonmagnetic classification accuracy, and the fraction of compounds with absolute prediction error below 0.1 $\mu_{\mathrm{B}}$ for all compounds and for the magnetic subset.
Since the goal of this study is to identify compounds with large $E_{\mathrm{aniso}}$, it is first necessary to determine whether a compound is magnetic. Relying solely on the total magnetization is inadequate, as it cannot capture antiferromagnetic (AFM) or low-moment ferrimagnetic (FiM) compounds. To address this, we employed the total absolute magnetic moment, defined as $\sum_{i}|m_{i}|$, where $m_{i}$ denotes the local magnetic moment at atomic site $i$.
Local magnetic moments were predicted using an MLRM based on the eSEN architecture, trained to output the moment at each atomic site. Restricting to collinear configurations in which all moments are aligned along the $z$-axis, each moment is represented by a scalar whose sign encodes the direction, with $\ell=0$ per site in the output head. To account for the $z$-direction ambiguity—where a configuration and its sign-inverted counterpart (i.e., all local moments flipped) are physically equivalent—we modified the loss function to compute losses for both the predicted $\{\boldsymbol{m}_{i}\}$ and its sign-inverted counterpart, and take the smaller value as the loss. This ensures invariance under global spin inversion.
Fig. 5 (a) shows a scatter plot comparing $\{\boldsymbol{m}_{i}\}$ from the MLRM and DFT for compounds in test set, with histograms along the axes illustrating their distributions. Local moments at the atom site $Z$ are omitted as they are nonmagnetic. Because 73.4% of the test compounds are nonmagnetic, both histograms exhibit a pronounced peak at zero. For magnetic systems, the global sign is adjusted so that the total magnetic moment is positive; since most magnetic compounds are ferromagnetic, positive moments dominate in the distribution. Nearly all points fall along the diagonal and only 1.4% of points lie along either axis, demonstrating that the model accurately predicts both the magnitude and sign of local moments, and reliably distinguishes ferromagnetic (FM) and ferrimagnetic (FiM) systems. For the magnetic/nonmagnetic classification, we evaluated performance using receiver operating characteristic (ROC) and precision-recall (PR) curves, as shown in Fig. S6. The area under the curve (AUC) values are 0.98 and 0.97, respectively, indicating highly accurate classification.
Predicting $\{\boldsymbol{m}_{i}\}$ is a common yet computationally demanding step in HTP studies of magnetic materials. The approach developed in this work achieves high accuracy in $\{\boldsymbol{m}_{i}\}$ prediction and is readily transferable to other systems. A central question, however, is how many training compounds are required to reach satisfactory accuracy. To address this, we performed a learning-curve analysis by training the MLRM on progressively larger subsets of the dataset and evaluating performance on a fixed test set of 5000 compounds, of which 1486 are magnetic. For evaluation, local moments at the $X_{1}$, $X_{2}$, and $Y$ sites were concatenated across samples into a single array, while the $Z$ site was excluded since it is nonmagnetic in conventional Heusler compounds.
The learning curve is shown in Fig. 5 (b), illustrating how model performance improves with increasing training set size. In the top panel, the $R^{2}$ scores for both local moments and their magnitudes are presented. The gap between the two curves indicates that, while the model generally captures the magnitude accurately, it sometimes assigns the incorrect sign. For example, Mn2ScGe with DFT-computed local moments {$2.62$, $3.03$, 0.29-0.29, 0.10-0.10} $\mu_{\mathrm{B}}$ is predicted as {2.71-2.71, $3.04$, $0.02$, 0.02-0.02} $\mu_{\mathrm{B}}$ by MLRM. The bottom panel reports two key metrics: (i) classification accuracy for identifying magnetic systems and (ii) the fraction of local moments with absolute error below 0.1 $\mu_{\mathrm{B}}$, evaluated across all compounds and within the magnetic subset. With 5000 training samples, the model achieves a classification accuracy of 0.92, and 90% of all compounds fall within the 0.1 $\mu_{\mathrm{B}}$ error threshold. However, this fraction decreases to 72% when restricted to magnetic compounds subset, indicating that the model identifies whether a site is magnetic with high reliability but remains less accurate in predicting exact $\{\boldsymbol{m}_{i}\}$ values. Increasing the training set to 125,000 samples improves this fraction to 82%, while relaxing the threshold to 0.2 $\mu_{\mathrm{B}}$ further raises it to 92%. Although performance improves with larger datasets, the gains become progressively smaller. These results highlight the critical role of dataset size in improving $\{\boldsymbol{m}_{i}\}$ accuracy and inform the selection of training size in future work focusing on other magnetic systems.